# PIKA: Expert-Level Synthetic Datasets for Post-Training Alignment from Scratch

### v1

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has become a cornerstone for aligning large language models (LLMs). However, its effectiveness critically depends on high-quality instruction data. Most existing high-quality alignment datasets are either private or require costly human annotation, which hinders reproducibility and scalability. Even with the emergence of Reinforcement Learning from AI Feedback (RLAIF), concerns about data quality remain. Moreover, it is still unclear to the open-source community how much data is actually required to fine-tune a base model into a strong instruction-following model. Current state-of-the-art approaches typically rely on over 300k examples even at the supervised fine-tuning (SFT) stage, yet they still underperform compared to proprietary models, leaving substantial barriers for academic and resource-constrained communities. To address this gap, we introduce PiKa, a data-efficient family of expert-level alignment datasets. In particular, the PiKa-SFT dataset uses only 30k SFT examples—an order of magnitude fewer than the SoTA dataset Magpie. Through extensive evaluations by fine-tuning Llama-3-8B-Base on PiKa and other public instruction following datasets, we show that PiKa-SFT alone outperforms models trained on much larger datasets. Remarkably, on two widely used alignment benchmarks, AlpacaEval 2.0 and Arena-Hard, PiKa-SFT fine-tuning surpasses the official Llama-3-8B-Instruct model – which was trained on over 10M proprietary examples. We further extend our study by training the Qwen2.5 series (0.5B–7B) on PiKa-SFT, consistently outperforming their official instruction-tuned counterparts. In addition, we curate 30k high-quality preference optimization examples, which further improve alignment performance when applied after SFT initialization. These findings demonstrate that high-quality alignment can be achieved with significantly reduced data, providing a practical and scalable path for advancing open-source LLM alignment research. Our code and data will be available at https://github.com/SJY8460/PiKa.
@@ 1 Introduction
To overcome the difficulty of building high-quality datasets, two major directions have been explored. The first relies on human experts to author and curate instruction–response pairs (Databricks, 2023; Köpf et al., 2023; Zhao et al., 2024; Zheng et al., 2024; 2023), a process that is both time-consuming and labor-intensive (Liu et al., 2024). The second line of work leverages LLMs themselves to automatically generate synthetic data (Ding et al., 2023; Yin et al., 2023; Li et al., 2024a; Sun et al., 2023; Taori et al., 2023; Wang et al., 2023; 2024; Xu et al., 2023a; b; Li et al., 2023a). More recently, efforts have shifted toward increasing the diversity of synthetic datasets and scaling them up to massive collections of conversational examples that aim to approximate human interaction. However, despite such progress, models trained from scratch on these large-scale synthetic corpora often acquire only limited instruction-following ability, and the training process itself requires significant computational resources, making it less accessible for much of the research community.
@@
Figure 1: Overview of the PiKa pipeline for synthesizing expert-level alignment data. Step 1: An expert-level persona covering multiple domains is given to an aligned LLM, which auto-regressively generates high-quality and knowledge-intensive instructions. Step 2: For each instruction, the LLM produces multiple candidate responses. Step 3: A reward model scores these responses. For SFT, we select the instruction and its highest-scoring response as a training pair. For DPO, we construct preference data by pairing the highest- and lowest-scoring responses with the same instruction.
@@
Building on these insights, we construct an expert-level synthetic instruction dataset named PiKa, as illustrated in Figure 1. Our approach is inspired by persona-based generation methods (Ge et al., 2025), which prompt LLMs to role-play as specific personas to generate diverse instructions. To further enhance data quality, we rigorously sample and filter persona prompts, removing harmful or policy-violating cases and retaining only complex personas for instruction generation. Unlike many existing datasets, PiKa is fully synthesized using GPT-4o, a model widely regarded as both highly capable and human-like, ensuring consistently high-quality outputs. Moreover, all generated instruction–response pairs are filtered and re-scored by GPT-4o, resulting in datasets dominated by challenging, expert-level prompts with detailed solutions. We provide comprehensive statistics, highlight the advantages of PiKa, and present an in-depth analysis in Section 3, enabling practitioners to flexibly select and filter data for fine-tuning according to their specific needs.
@@
To evaluate the effectiveness of PiKa, we compare it with widely used public instruction datasets (e.g., ShareGPT (Chiang et al., 2023), WildChat (Zhao et al., 2024), UltraChat (Ding et al., 2023), OpenHermes (Teknium, 2023a; b), Tulu V2 Mix (Ivison et al., 2023), and Magpie-Air/Pro (Xu et al., 2025)) by conducting supervised fine-tuning (SFT) of the Llama-3-8B-Base model on each dataset and evaluating the resulting models on standard alignment benchmarks such as AlpacaEval 2 (Li et al., 2023b) and Arena-Hard (Li et al., 2024b), following recent Magpie studies. Our results show that models fine-tuned on PiKa consistently achieve superior performance, even surpassing those trained with both SFT and direct preference optimization (DPO) (Rafailov et al., 2023) on UltraFeedback (Cui et al., 2023). Remarkably, PiKa is more than an order of magnitude smaller than Magpie-Pro (30K vs. 300K examples for SFT). Despite this significant difference in dataset size, PiKa achieves superior performance across benchmarks. For the first time, our approach enables open-source aligned models to significantly outperform the official Llama-3-8B-Instruct on both AlpacaEval 2.0 and Arena-Hard. This is particularly remarkable given that Llama-3-8B-Instruct was trained with over 10M proprietary examples and subsequent preference optimization. Building on this, we extend our study to the Qwen2.5 (0.5B–7B), where models trained on PiKa-SFT consistently surpass their official instruction-tuned counterparts.
@@
Furthermore, PiKa not only excels under SFT alone compared to prior public datasets, but also delivers state-of-the-art results when combined with preference optimization methods such as DPO. By extending PiKa to generate high-quality preference optimization data, PiKa-aligned Llama-3 models achieve substantial improvements and outperform models trained with both the Magpie and UltraFeedback series, with particularly strong gains on the more challenging Arena-Hard benchmark (43.70 vs. 33.30 for Magpie-Pro and 25.30 for UltraFeedback respectively). These findings highlight the exceptional quality and efficiency of PiKa, demonstrating that carefully constructed synthetic data can outperform extensively optimized proprietary datasets in a far more data-efficient manner.
@@ 2 PiKa: Expert-Level Synthetic Datasets
Step 1: Expert-Level Instruction Generation. We sample a diverse set of complex personas $\{\pi_{i}\}_{i=1}^{N}$ from PersonaHub (Ge et al., 2025), covering multiple domains such as biology, engineering, medicine, and law. Each persona $\pi_{i}$ is then provided to GPT-4o, which auto-regressively generates a corresponding knowledge-intensive instruction. Based on our preliminary experiments, we find that more domain-specific and sophisticated personas tend to generate more challenging prompts:
@@
Step 3: Reward-Model-Guided Selection. For each instruction $I_{i}$ ($i=1,\dots,N$), a reward model $\mathcal{R}(\cdot)$ assigns a score to each candidate responses. In our default setting, we employ Skywork-Reward-V2-Llama-3.1-8B (Liu et al., 2025), which ranks first on the RewardBench leaderboard [^1] https://huggingface.co/spaces/allenai/reward-bench, as of 2025/9/25. The reward score for each candidate response $r_{ij}$ of instruction $I_{i}$ is computed as:
@@
For supervised fine-tuning (SFT), we retain the pair $(I_{i},r_{ij^{*}})$ with
@@ 3 Dataset Analysis
Figure 2: Comparison between PiKa and Magpie-Pro. (a) Minimum input embedding distances distribution. (b) Instruction lengths distribution. (c) Response lengths Distribution.
@@ 3.1 Instruction Similarity
where $e(I)$ denotes the embedding of instruction $I$. Larger MND indicates lower redundancy and higher diversity. As shown in Figure 2(a), PiKa achieves a mean MND of 0.497 compared to Magpie-Pro’s 0.598, indicating comparable diversity while effectively avoiding repetitive prompts that plague many existing datasets.
@@ 3.2 Length Analysis
We further analyze tokenized lengths of instructions and responses. Figure 2(b, c) shows that PiKa instructions average around 424 tokens, and responses exceed 5,305 tokens on average—substantially longer than existing datasets—indicating that PiKa provides more detailed and knowledge-intensive supervision.
@@ 3.3 Dataset Assessment
Beyond coverage and length, we adopt GPT-4o as an expert evaluator to assess difficulty, feasibility, and quality . Each attribute is scored on a 1 to 10 scale with dedicated prompts (see Appendix C). Formally, let
$s_{d}$ $:\mathcal{X}\to[1,10],$ (7)
$s_{f}$ $:\mathcal{X}\to[1,10],$ (8)
$s_{q}$ $:\mathcal{X}\times\mathcal{Y}\to[1,10],$ (9)
denote the difficulty of instruction $x$, the feasibility of instruction $x$, and the quality of instruction response pair $(x,y)$, respectively.
Figure 3: GPT-4o–based evaluation of PiKa’s instruction difficulty, feasibility and instruction-response quality compared with Magpie-Pro.
@@ 3.3.1 Difficulty Evaluation
Figure 4: Example instructions from Magpie-Pro and PiKa datasets. PiKa instructions are generally more complex and domain-specific, while Magpie-Pro tends to focus on simpler, open-ended prompts. We also include detailed instruction-response pairs in Appendix D.
@@ 3.3.3 Quality Evaluation
Table 1: Performance comparison of instruction-tuned models based on Llama-3-8B using PiKa-generated versus baseline datasets. PiKa achieves superior performance while requiring $10\times$ less training data than state-of-the-art Magpie methods. Standard deviations (SD) reported are computed across both LC and WR metrics and #Convs indicates the number examples used for training the base model.
Alignment Setup
(Base LLM = Llama-3-8B)
#Convs
AlpacaEval 2 Arena-Hard
@@ 4.1 Experimental Setup
Baselines for Supervised Fine-Tuning and Preference Optimization. We compare PiKa with nine state-of-the-art open-source instruction datasets: Self-Instruct (Wang et al., 2023), ShareGPT (Chiang et al., 2023), WildChat (Zhao et al., 2024), UltraChat (Ding et al., 2023), OpenHermes 1 (Teknium, 2023a), OpenHermes 2.5 (Teknium, 2023b), Tulu V2 Mix (Ivison et al., 2023), Magpie-Air (Xu et al., 2025), and Magpie-Pro (Xu et al., 2025). ShareGPT and WildChat represent human-authored datasets containing 112K and 652K high-quality multi-turn conversations between humans and ChatGPT, respectively. UltraChat and the Magpie family are representative open-source synthetic datasets. For preference optimization, we compare models aligned using PiKa with direct preference optimization (DPO) (Rafailov et al., 2023) baselines, specifically comparing against two state-of-the-art open-source training datasets: UltraFeedback and Magpie-Pro.
@@
Evaluation Benchmarks. We evaluate aligned model performance using two widely adopted instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023b) and Arena-Hard (Li et al., 2024b). AlpacaEval 2 comprises 805 representative instructions selected from real user interactions, providing a comprehensive assessment of practical instruction-following capabilities. Arena-Hard represents an enhanced version of MT-Bench (Zheng et al., 2023), containing 500 challenging user queries designed to test advanced reasoning and problem-solving abilities. Both benchmarks employ GPT-based evaluators to assess responses generated by the model under evaluation against established baseline models. Specifically, we use GPT-4-Turbo (1106) and Llama-3-8B-Instruct as baselines for AlpacaEval 2 and Arena-Hard uses GPT-4 (0314) as its baseline. In addition, we employ the latest GPT-4o as the LLM-as-judge to provide more contemporary and robust evaluation standards.
@@
4.2 Experimental Results
[…]
Figure 5: Comparison across three dataset families: UltraFeedback, Magpie-Pro, and PiKa. (a) Dataset sizes. (b) AlpacaEval 2.0 win rates. (c) Arena-Hard win rates. .
Performance on Preference Optimization. To evaluate PiKa’s effectiveness in preference optimization, we compare it against UltraFeedback and Magpie-Pro across data efficiency and benchmark performance (Figure 5). As shown in Figure 5(a), PiKa is highly data-efficient, requiring only 30K examples for SFT and 30K for DPO. In contrast, UltraFeedback uses 208K and 64K, while Magpie-Pro uses 300K and 60K, respectively. Despite this substantial reduction in training data, PiKa achieves superior results on both evaluation benchmarks. On AlpacaEval 2.0, PiKa attains a 41.29% win rate, significantly surpassing Magpie-Pro (36.24%) and UltraFeedback (18.40%). Similarly, on Arena-Hard, PiKa achieves 43.70%, outperforming Magpie-Pro (33.30%) and UltraFeedback (25.30%). These results highlight PiKa’s ability to generate high-quality preference data, enabling more effective alignment with human preferences even under constrained data budgets.
(a) AlpacaEval 2.0 LC scaling
(b) AlpacaEval 2.0 WR scaling
(c) Arena-Hard WR scaling
Table 3: Performance comparison on additional benchmarks from the Open LLM Leaderboard. All models are supervised fine-tuned on Llama-3-8B base models. Numbers in parentheses indicate the number of few-shot examples used for each task.

### v2

@@ Abstract
High-quality instruction data is critical for LLM alignment, yet existing open-source datasets often lack efficiency, requiring hundreds of thousands of examples to approach proprietary performance. In this work, we find that beyond the widely recognized importance of prompt–response quality, prompt difficulty itself plays a critical role in driving alignment gains. Motivated by this observation, we introduce PiKa, a data-efficient family of expert-level alignment datasets that concentrates supervision on high-difficulty instructions. The PiKa-SFT dataset contains only 30k examples—an order of magnitude fewer than state-of-the-art open datasets like Magpie-Pro. Despite its small size, fine-tuning Llama-3-8B-Base on PiKa-SFT even outperforms the official Llama-3-8B-Instruct model—trained on over 10M proprietary examples—on widely used benchmarks such as AlpacaEval 2.0 and Arena-Hard. We also validate the generalizability of PiKa across the Qwen2.5 series (0.5B–7B), consistently surpassing their official instruction-tuned counterparts. Additionally, we provide 30k high-quality preference optimization examples to further enhance alignment. Our results demonstrate that promising alignment is achievable with significantly reduced data, democratizing access for resource-constrained research. Our code and data will be available at https://github.com/SJY8460/PiKa.
PiKa: Expert-Level Synthetic Datasets for Post-Training Alignment from Scratch
Shangjian Yin1,2   Shining Liang2   Wenbiao Ding2   Yuli Qian2 Zhouxing Shi1   Hongzhi Li2   Yutao Xie2 1University of California, Riverside   2Microsoft AI
@@ 1 Introduction
To overcome the difficulty of building high-quality datasets, two major directions have been explored. The first relies on human experts to author and curate instruction–response pairs (Databricks, 2023; Köpf et al., 2023; Zhao et al., 2024; Zheng et al., 2024, 2023), a process that is both time-consuming and labor-intensive (Liu et al., 2024). The second line of work leverages LLMs themselves to automatically generate synthetic data (Ding et al., 2023; Yin et al., 2023; Li et al., 2024a; Sun et al., 2023; Taori et al., 2023; Wang et al., 2023, 2024; Xu et al., 2023a, b; Li et al., 2023a). More recently, efforts have shifted toward increasing the diversity of synthetic datasets and scaling them up to massive collections of conversational examples that aim to approximate human interaction. However, despite such progress, models trained from scratch on these large-scale synthetic corpora often acquire only limited instruction-following ability, and the training process itself requires significant computational resources, making it less accessible for much of the research community.
@@
Figure 1: Overview of the PiKa pipeline for synthesizing expert-level alignment data.
@@
Building on these insights, we construct an expert-level synthetic instruction dataset named PiKa, as illustrated in Figure 1. Our approach is inspired by persona-based generation methods (Ge et al., 2025), which prompt LLMs to role-play as specific personas to generate diverse instructions. To further enhance data quality, we rigorously sample and filter persona prompts, removing harmful or policy-violating cases and retaining only complex personas for instruction generation. Unlike many existing datasets, PiKa is fully synthesized through an automated generation pipeline, enabling scalable and consistent data construction. All generated instruction–response pairs are subsequently filtered and re-scored to prioritize difficulty and solution quality, resulting in a dataset dominated by challenging, expert-level prompts paired with detailed solutions. We provide comprehensive statistics, highlight the advantages of PiKa, and present an in-depth analysis in Section 3, enabling practitioners to flexibly select and filter data for fine-tuning according to their specific needs.
@@
To evaluate the effectiveness of PiKa, we compare it with widely used public instruction datasets (e.g., ShareGPT (Chiang et al., 2023), WildChat (Zhao et al., 2024), UltraChat (Ding et al., 2023), OpenHermes (Teknium, 2023b, a), Tulu V2 Mix (Ivison et al., 2023), and Magpie-Air/Pro (Xu et al., 2025)) by conducting supervised fine-tuning (SFT) of the Llama-3-8B-Base model on each dataset and evaluating the resulting models on standard alignment benchmarks such as AlpacaEval 2 (Li et al., 2023b) and Arena-Hard (Li et al., 2024b), following recent Magpie studies. Our results show that models fine-tuned on PiKa consistently achieve superior performance, even surpassing those trained with both SFT and direct preference optimization (DPO) (Rafailov et al., 2023) on UltraFeedback (Cui et al., 2023). Remarkably, PiKa is more than an order of magnitude smaller than Magpie-Pro (30K vs. 300K examples for SFT). Despite this significant difference in dataset size, PiKa achieves superior performance across benchmarks. For the first time, our approach enables open-source aligned models to significantly outperform the official Llama-3-8B-Instruct on both AlpacaEval 2.0 and Arena-Hard. This is particularly remarkable given that Llama-3-8B-Instruct was trained with over 10M proprietary examples and subsequent preference optimization. Building on this, we extend our study to the Qwen2.5 (0.5B–7B), where models trained on PiKa-SFT consistently surpass their officially tuned counterparts.
@@
Furthermore, PiKa not only excels under SFT alone compared to prior public datasets, but also delivers the best results when combined with preference optimization methods such as DPO. By extending PiKa to generate high-quality preference optimization data, PiKa-aligned Llama-3 models achieve substantial improvements and outperform models trained with both the Magpie and UltraFeedback series, with particularly strong gains on the more challenging Arena-Hard benchmark (43.70 vs. 33.30 for Magpie-Pro and 25.30 for UltraFeedback respectively). These findings highlight the exceptional quality and efficiency of PiKa, demonstrating that synthetic data constructed by prioritizing prompt difficulty and expert-level supervision can outperform extensively optimized proprietary datasets in a far more data-efficient manner.
@@ 2 PiKa: Expert-Level Synthetic Datasets
Our primary contribution is the construction of PiKa, a family of expert-level synthetic datasets for data-efficient alignment. This section introduces the design principles and generation framework underlying the PiKa datasets.
Step 1: Expert-Level Instruction Generation. We sample a diverse set of personas $\{\pi_{i}\}_{i=1}^{N}$ from PersonaHub (Ge et al., 2025), covering multiple domains such as biology, engineering, medicine, and law. Each persona $\pi_{i}$ is then provided to GPT-4o, which auto-regressively generates a corresponding knowledge-intensive instruction. Based on our preliminary experiments, we find that more domain-specific and sophisticated personas tend to generate more challenging prompts:
@@
Step 3: Reward-Model-Guided Selection. For each instruction $I_{i}$ ($i=1,\dots,N$), a reward model $\mathcal{R}(\cdot)$ assigns a score to each candidate responses. In our default setting, we employ Skywork-Reward-V2-Llama-3.1-8B (Liu et al., 2025), which ranks first on the RewardBench leaderboard [^1] https://huggingface.co/spaces/allenai/reward-bench, as of 2026/4/8. The reward score for each candidate response $r_{ij}$ of instruction $I_{i}$ is computed as:
@@
For SFT, we retain the pair $(I_{i},r_{ij^{*}})$ with
@@ 3 Dataset Analysis
Figure 2: (a) Minimum input embedding distances distribution. (b) Instruction lengths distribution. (c) Response lengths Distribution.
@@ 3.1 Instruction Similarity
where $e(I)$ denotes the embedding of instruction II. Larger MND indicates lower redundancy and higher diversity. As shown in Figure 2(a), PiKa achieves a mean MND of 0.497 compared to Magpie-Pro’s 0.598, indicating comparable diversity while effectively avoiding repetitive prompts that plague many existing datasets.
@@ 3.2 Length Analysis
We further analyze the tokenized lengths of instructions and responses. Figure 2(b, c) shows that PiKa instructions average around 424 characters, and responses exceed 5,305 characters (around 1300 tokens) on average—substantially longer than existing datasets—indicating that PiKa provides more detailed and knowledge-intensive supervision.
@@ 3.3 Dataset Assessment
Beyond coverage and length, we adopt GPT-4o as an expert evaluator to assess difficulty, feasibility, and quality. Each attribute is scored on a 1–10 scale using dedicated prompts (see Appendix B). Formally, we define $s_{d}:\mathcal{X}\!\to\![1,10]$, $s_{f}:\mathcal{X}\!\to\![1,10]$, and $s_{q}:\mathcal{X}\!\times\!\mathcal{Y}\!\to\![1,10]$, which denote the difficulty of instruction $x$, the feasibility of instruction $x$, and the quality of an instruction–response pair $(x,y)$, respectively.
Figure 3: LLM–based evaluation of PiKa’s instruction difficulty, feasibility and instruction-response quality compared with Magpie-Pro. We also provide results from multiple judges in Section E.7.
@@ 3.3.3 Quality Evaluation
Table 1: Performance comparison of instruction-tuned models based on Llama-3-8B-Base using PiKa-generated versus baseline datasets. PiKa achieves superior performance while requiring $10\times$ less training data than state-of-the-art Magpie methods. Standard deviations (SD) reported are computed across both LC and WR metrics and #Convs indicates the number examples used for training the base model.
Alignment Setup (Base LLM = Llama-3-8B-Base) #Convs AlpacaEval 2 Arena-Hard
@@ 4.1 Experimental Setup
Baselines for Supervised Fine-Tuning and Preference Optimization. We compare PiKa with nine state-of-the-art open-source instruction datasets: Self-Instruct (Wang et al., 2023), ShareGPT (Chiang et al., 2023), WildChat (Zhao et al., 2024), UltraChat (Ding et al., 2023), OpenHermes 1 (Teknium, 2023b), OpenHermes 2.5 (Teknium, 2023a), Tulu V2 Mix (Ivison et al., 2023), Magpie-Air (Xu et al., 2025), and Magpie-Pro (Xu et al., 2025). ShareGPT and WildChat represent human-authored datasets containing 112K and 652K high-quality multi-turn conversations between humans and ChatGPT, respectively. UltraChat and the Magpie family are representative open-source synthetic datasets. For preference optimization, we compare models aligned using PiKa with direct preference optimization (DPO) (Rafailov et al., 2023) baselines, specifically comparing against two state-of-the-art open-source training datasets: UltraFeedback and Magpie-Pro.
@@
Evaluation Benchmarks. We evaluate aligned model performance using two widely adopted instruction-following benchmarks: AlpacaEval 2 (Li et al., 2023b) and Arena-Hard (Li et al., 2024b). AlpacaEval 2 comprises 805 representative instructions selected from real user interactions, providing a comprehensive assessment of practical instruction-following capabilities. Arena-Hard represents an enhanced version of MT-Bench (Zheng et al., 2023), containing 500 challenging user queries designed to test advanced reasoning and problem-solving abilities. Both benchmarks employ GPT-based evaluators to assess responses generated by the model under evaluation against established baseline models. Specifically, we use GPT-4-Turbo (1106) and Llama-3-8B-Instruct as baselines for AlpacaEval 2 and Arena-Hard uses GPT-4 (0314) as its baseline. In addition, we primarily employ the latest GPT-4o as the LLM-as-judge to provide contemporary and robust evaluation standards. We also include evaluations using GPT-5 and GPT-4.1-mini, as shown in Table 4.
@@
Table 3: Ablation study on prompt difficulty. We compare the Magpie-Pro baseline with different PiKa variants. We also provide an ablation on the generator choice for PiKa in Section E.9.
AlpacaEval 2 Dataset Difficulty LC (%) WR (%) Magpie-Pro 2.65 15.42 16.89 PiKa-Series (10k Subset)      w/o Persona-Guide 3.11 13.84 15.53      Low-Diff 2.91 21.86 14.95      Mid-Diff 3.64 24.36 17.84      Expert (Default) 7.39 31.01 30.32
Table 4: Additional evaluation results on AlpacaEval 2 judged by GPT-5 and GPT-4.1-mini. Note: LC is the most important metric.
AlpacaEval 2 Judger Model LC (%) WR (%) GPT-5 Official-Instruct 15.95 16.56 Magpie-Pro 9.51 14.55 PiKa 19.28 19.00 GPT-4.1-mini Official-Instruct 25.44 25.53 Magpie-Pro 20.17 28.54 PiKa 29.02 27.55
Table 5: Performance comparison on additional downstream objective tasks from the Open LLM Leaderboard. The goal of this evaluation is to assess whether alignment with PiKa preserves performance on objective tasks, rather than to optimize for further gains. All models are supervised fine-tuned on Llama-3-8B-Base. Numbers in parentheses indicate the number of few-shot examples.
@@
4.2 Experimental Results
[…]
Figure 4: Comparison across three dataset families: UltraFeedback, Magpie-Pro, and PiKa. .
(a) AlpacaEval 2.0 LC scaling
(b) AlpacaEval 2.0 WR scaling
(c) Arena-Hard WR scaling
Expert-level prompt difficulty is a key driver of performance gains. To verify the effectiveness of the difficulty-driven construction strategy, we investigate the impact of prompt difficulty on the model’s generative chat ability. We trained models on 10k subsets across different categories. As shown in Table 3, the w/o Persona-Guide variant fails to surpass the baseline, indicating that difficulty alone is insufficient without persona constraints. Crucially, we observe a positive correlation between prompt difficulty and chat performance. As the difficulty score rises from 2.91 (Low-Diff) to 3.64 (Mid-Diff) and finally to 7.39 (Expert), the generative chat ability improves significantly. Specifically, the Expert subset boosts LC from 21.86% to 31.01% and doubles the WR from 14.95% to 30.32%, suggesting that prompt difficulty is a key driver for superior chat capabilities.
Preference optimization further amplifies performance gains. To evaluate PiKa’s effectiveness in preference optimization, we compare it against UltraFeedback and Magpie-Pro across data efficiency and benchmark performance (Figure 4). As shown in Figure 4(a), PiKa is highly data-efficient, requiring only 30K examples for SFT and 30K for DPO. In contrast, UltraFeedback uses 208K and 64K, while Magpie-Pro uses 300K and 60K, respectively. Despite this substantial reduction in training data, PiKa achieves superior results on both evaluation benchmarks. On AlpacaEval 2.0, PiKa attains a 41.29% win rate, significantly surpassing Magpie-Pro (36.24%) and UltraFeedback (18.40%). Similarly, on Arena-Hard, PiKa achieves 43.70%, outperforming Magpie-Pro (33.30%) and UltraFeedback (25.30%). These results highlight PiKa’s ability to generate high-quality preference data, enabling more effective alignment with human preferences even under constrained data budgets.